The perimeter of a rectangle is 54 meters. The longer side is 3 meters longer than the shorter side. Find the length of each side.
Answer by longjonsilver(2297)

Perimeter is "add up all the lengths"

Let short side be x, so the longer side is? well 3m longer, so it is x+3.

So, a rectangle has 4 sides...2 of one length and the other 2 of the other length:

x + x + x+3 + x+3 = 54
4x + 6 = 54
4x = 48
x = 48/4
--> x = 12m

So, short side is 12m, therefore longer side must be 15m.

CHECK: 12+12+15+15 is 54... correct answer.

jon.